NOAA SATOPS Morning Report: Friday, November 02, 2012


SATOPS Morning Report: Friday, November 02, 2012

Customers or users who want to call to ask about problems or report outages can use the following contacts:
- For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444
- To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: LRIT@noaa.gov

DOY 307

GOES
GOES-14 DOY305 19:00z stopped imaging and sounding schedules per plan.
GOES-14 will remain on standby at 89.5 degrees West until further notice.

All other GOES operations were nominal over the past 24 hours.
No anomalies affecting product delivery via GVAR or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
N17 DOY307 21:00z output power of the transmitter used for HRPT data dropped to red low limits. Problem under investigation.

All other POES operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
All DMSP operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
All JASON-2 operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.

ESPC Reported data outage and delays
Monterey reported to ESPC that DMSP F17 Rev. 30928 will be lost due to problems ingesting this pass.

Questions or comments to Ron.Mahmot@noaa.gov

Please follow SpaceRef on Twitter and Like us on Facebook.